OpenStudy (beth12345):

Waves travel 0.75 times as fast in shallow water as they do in deep water. What will be the wavelength of the waves in deep water, if their wavelength is 2.0cm in shallow water?

OpenStudy (beth12345):

the answer im given is 2.7cm, i just don't know how to get it, if someone could explain, i would be grateful :)

OpenStudy (anonymous):

velocity=frequency*wavelength. constant frequency. v2/v1=.75=2*f/(x*f) solve for x

OpenStudy (anonymous):

velocity = frequency x wavelength v1= f x WL1 frequency doesn't change. v2 = f x WL2 for v2 to be v1/0.75 = > WL2 must be WL1/0.75= 2/0.75=2.6666 cm